OpenStudy (anonymous):

Simplify this expression: 4p + 9 + (–7p) + 2

OpenStudy (haseeb96):

4p-7p +9 +2 =0 -3p + 11=0 11=3p p= 11/3 this is the correct answer

OpenStudy (anonymous):

=4p+9-7p+2 =(4-7)p+11 =-5p+11

OpenStudy (haseeb96):

how old are u ? @kmckn

OpenStudy (anonymous):

oh sorrey...-3p+11 not -5p+11

OpenStudy (anonymous):

@Haseeb96 its not equal to zero..it is just to simplify..